Tohban Report 2011-10-12

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 5 Oct 2011
End Date: 12 Oct 2011
Tohban: Hazel Bain
Tohban email:
Next Tohban: Jim McTiernan
List all reports

Solar Activity

Solar activity was low with only C class flares. There are currently nine active regions on disk, two of which will rotate over the limb within the next 24 hours. STEREO B shows one further active region due to rotate onto the disk within the next 48 hours. Due to problems with RHESSI detector 5 (see below) which caused the SSR level to reach 95% we were forced to delete data from 6-oct-2011 04:20 to 11:55, causing us to miss two C class flares.

How many GOES flares occurred?

 Flares above B, C, M, X class were     18    18     0     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      5     6     0     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      3     2     0     0

There were RHESSI flares/GOES flares 100 / 36 over the time range 05-Oct-11 12-Oct-11


At 01:30 on 6 Oct 2011 detector 5 showed a sudden increase of low energy (a few keV) counts from the front segment.

The cause of this is still unknown. The increase in low energy counts caused the SSR to gradually fill, reaching a max of around 95%. During the second BGS pass on the 7th Oct 2011 (2011-280-02:58 UT) the SSR pointer was skipped to lower the SSR fill level to about 50%, deleting around 8 hours of data from 6-oct-2011 04:20 to 11:55. Detector 5's voltage was also dropped from 170 to 130. However the counts in detector 5 remained high. On the third pass the front segment fast and slow event thresholds were raised to their maximum, "0xFF." This was enough to lower the count rate. During this pass detector 5's voltage was then raised up to 150. On the 8th Oct 2011 The front segment fast and slow event thresholds lowered to "0x33." and "0x30", removing the low energy counts from detector 5. However since the pass was only 5 minutes long there wasn't time for further tweaking to optimize the thresholds.

A spike in the detector 5 monitor rates was observed on the 8th Oct 2011 prompting us to change the combination of detectors used to average livetime calculation used to trigger the attenuators from 1, 2, 3 and 5 to 1, 2, 3 and 4.

A few isolated spikes were observed in detector 8 followed by a longer periods of noise on the 11th of October. If this continues, the HV will be lowered.

Data Gaps

GAP START TIME              GAP END TIME                   GAP (SEC)

2011-10-06T04:20:00.000 -- 2011-10-06T11:55:00.000 27300.000

Personal tools